How much thermal energy must be transferred to a 20.0-g piece of copper wire to raise it from room temperature (25.0°C) to its melting temperature (1085.0°

CCu = 385 J/(kg·°C)
1
Expert's answer
2020-05-24T17:58:59-0400

We can find the quantity of the thermal energy from the formula:


"Q = mc \Delta T,"

here, "Q" is the quantity of the thermal energy that must be transferred to a 20.0-g piece of copper wire to raise it from room temperature (25.0°C) to its melting temperature (1085.0°C), "m = 0.02 kg" is the mass of the piece of copper wire, "c = 385 \dfrac{J}{kg \cdot \! ^{\circ}C}" is the specific heat capacity of the copper, "\Delta T" is the change in temperature.

Then, we get:


"Q = 0.02kg \cdot 385 \dfrac{J}{kg \cdot \! ^{\circ}C} \cdot (1085 \ ^{\circ}C - 25 \ ^{\circ}C) = 8162J."

Answer:

"Q = 8162J."